OpenStudy (anonymous):

Two charges separated a distance of 1.0 meter exert a 2.0-N force on each other. If the charges are pushed to a separation of 1/3 meter, the force on each charge will be A. 2.0 N. B. 1.5 N. C. 3.0 N. D. 6.0 N. E. 18.0 N.

OpenStudy (nayef):

\[F=k\frac{ Q1*Q2 }{ d ^{2}}\] d=seperation If we suppose the k*Q1*Q2=N ,where is just a represntation symbol then \[F=\frac{ N }{ d ^{2} }\] and \[\frac{ N }{ 1^{2} }=2 , then N=2 \] \[\frac{ 2 }{ (\frac{ 1 }{ 3 })^{2} }=18\] answer 18

OpenStudy (nayef):

where N is just a represntation symbol.

OpenStudy (anonymous):

Thank you :)